What is part d) the DOL? A project has the following estimated data: Price = $60 per unit; variable costs = $37 per unit; fixed costs = $21,500; requi
Answer #1

Degree of operating leverage = 1+(fixed cost/OCF)

OCF= investment/ PVIFA 12%, 3 years ( 3 year annuity Factor at 12%)

Annuity Factor = 1 -(1+r)^ -n /r

Annuity Factor= 1-(1+.12)^-3 /.12= 0.28822/.12= 2.40183

3 year annuity Factor at 12%= 2.40183

OCF=18000/2.40183= 7494.28560

DOL= 1+(21500/7494.28560)=3 869

Note: as per your question only solved DOL.